Carter Company sells $500,000 of 10% bonds on November 1, 2018. The bonds yield 12% and pay interest on May 1 and November 1. The bonds are due on May 1, 2022. Bond premium interest or discount is amortized using the effective interest method, and is recorded in financial statements at interest dates and year-end. The year-end for Carter Company is December 31.

Required:

  1. Calculate the price of the bonds on the issuance date.
  2. Create an amortization schedule in Excel for the bonds.
  3. Prepare all relevant journal entries for this bond issue from the date of issuance through May 2020.
Answer #1

a. Maturity amount of the bonds = $ 500,000

Semi-annual coupons = $ 500,000 x 10 % x 1/2 = $ 25,000.

n = 7

i = 12 % x 1/2 = 6 %

PVA 6 %, n=7 = [ { 1 - ( 1 / 1.06 ) 7 } / 0.06 ] = 5.5824

PV 6 %, n = 7 = ( 1 / 1.06 ) 7 = 0.6651

Price of the bonds = 25,000 x 5.5824 + 500,000 x 0.6651 = 139,560 + 332,550 = $ 472,110

b. Amortization Table:

Date Amount Paid Interest Expense Discount Amortization Unamortized Discount Carrying Value of Bonds
Nov 1, 2018 $ 27,890 $ 472,110
May 1, 2019 $ 25,000 $ 28,326.60 $ 3,326.60 24,563.40 475,436.60
Nov 1, 2019 25,000 28,526.20 3,526.20 21,037.20 478,962.80
May 1, 2020 25,000 28,737.77 3,737.77 17,299.43 482,700.57
Nov 1, 2020 25,000 28,962.03 3,962.03 13,337.40 486,662.60
May 1, 2021 25,000 29,199.76 4,199.76 9,137.64 490,862.36
Nov 1, 2021 25,000 29,451.74 4,451.74 4,685.90 495,314.10
May 1, 2022 25,000 29,685.90 4,685.90 0 500,000

c. In the books of Carter Com[pany :

Date Account Titles Debit Credit
$ $
Nov 1, 2018 Cash 472,110
Discount on Bonds Payable 27,890
Bonds Payable 500,000
May 1, 2019 Interest Expense 28,326.60
Discount on Bonds Payable 3,326.60
Cash 25,000
Nov 1, 2019 Interest Expense 28,526.20
Discount on Bonds Payable 3,526.20
Cash 25,000
Dec 31, 2019 Interest Expense 9,579.25
Discount on Bonds Payable 1,245.92
Interest Payable 8,333.33
May 1, 2020 Interest Expense 19,158.52
Interest Payable 8,333.33
Discount on Bonds Payable 2,491.85
Cash 25,000